# System Identification of Neural Systems: Going Beyond Images to Modelling Dynamics

### v1

@@ Abstract
Vast literature has compared the recordings of biological neurons in the brain to deep neural networks. The ultimate goal is to interpret deep networks or to better understand and encode biological neural systems. Recently, there has been a debate on whether system identification is possible and how much it can tell us about the brain computation. System identification recognizes whether one model is more valid to represent the brain computation over another. Nonetheless, previous work did not consider the time aspect and how video and dynamics (e.g., motion) modelling in deep networks relate to these biological neural systems within a large-scale comparison. Towards this end, we propose a system identification study focused on comparing single image vs. video understanding models with respect to the visual cortex recordings. Our study encompasses two sets of experiments; a real environment setup and a simulated environment setup. The study also encompasses more than 30 models and, unlike prior works, we focus on convolutional vs. transformer-based, single vs. two-stream, and fully vs. self-supervised video understanding models. The goal is to capture a greater variety of architectures that model dynamics. As such, this signifies the first large-scale study of video understanding models from a neuroscience perspective. Our results in the simulated experiments, show that system identification can be attained to a certain level in differentiating image vs. video understanding models. Moreover, we provide key insights on how video understanding models predict visual cortex responses; showing video understanding better than image understanding models, convolutional models are better in the early-mid regions than transformer based except for multiscale transformers that are still good in predicting these regions, and that two-stream models are better than single stream.
@@ 1 Introduction
There has been a recent increase in studies that compare how deep neural networks process input stimuli to the processing that occurs in the brain whether in humans Zhou et al. (2022); Conwell et al. (2021b); Schrimpf et al. (2018); Cichy et al. (2019, 2021), non-human primates, or rodents Conwell et al. (2021a); Schrimpf et al. (2018). The benefits of these studies are two-fold. First, such a comparison can be used to interpret and have a better understanding of black-box deep neural networks and even provide inspirations on how to improve them. Second, it can provide a better understanding and encoding of biological neural systems which is the focus of this work. Towards achieving the latter, recent benchmarks have been released to improve the capabilities of machine learning models in neural encoding and comparing them to biological neural systems Schrimpf et al. (2018); Cichy et al. (2019, 2021); Gifford et al. (2023). The current established benchmark from The Mini-Algonauts Project 2021 Cichy et al. (2021), has provided neuro-imaging data for the brain responses from participants watching short video clips. Such a benchmark enables neuroscientists to study the question of how the brain understands actions and dynamics (e.g., motion).
Machine learning models have been investigated in encoding such neuro-imaging data Lahner (2022); Zhou et al. (2022). One recent approach proposed an improved regression scheme that considers the hierarchical nature of deep networks  Zhou et al. (2022). The method does not assume a one-to-one mapping between the layers and the regions, but rather learns a re-weighted combination of all selected layers in a deep network mapped to different visual regions. Yet, the focus of that study was on single image architectures that are not designed for video understanding, except for one model (i.e., TimeSformer) that was included in their study. Thus, there is still a lack of a unified large-scale study of deep video understanding models. Such a study can help us compare the dynamics modelling and action understanding between deep spatiotemporal networks and biological neural systems. Moreover, the previous study did not include any video understanding models that are trained in a self-supervised setting. There is an increased interest in self-supervised training, due to the rise of foundation models Bommasani et al. (2021). Foundation models refer to the set of models trained on broad datasets with tasks that allow powerful generalization typically in a self-supervised manner.
However, there is a question that needs to be investigated first before tackling such comparison. Is system identification feasible in both the real environment with the target model as the human visual cortex and a simulated environment where deep networks are used as the target model? A recent work has started such a quest to verify whether system identification is plausible or not Han et al. (2023). Yet, they focused on single image architectures, and their real environment data was a simple set of data from Brain-Score dataset Schrimpf et al. (2018) that does not include video stimuli. Additionally, their study used a simple encoding model for deep regression on brain signals that did not take the hierarchical nature and complex interactions between layers into consideration.
In this paper, our goal is to answer this system identification question and provide a large-scale study of video understanding models in encoding the human visual cortex. Furthermore, we aim to address the previously mentioned gaps in the literature. Our contributions are threefold:
@@
  • We conduct a system identification study that goes beyond single images to videos. We establish the simulated environment setup using video understanding models as the target and the real environment setup with the target model as the human visual cortex.
  • We showcase the first large-scale study of deep video understanding models including convolutional vs. transformer-based, single vs. two stream and fully vs. self-supervised.
  • We improve the encoding model in system identification, where we account for the hierarchical nature and complex cross-scale interactions through the previously proposed weighted neural encoding scheme Zhou et al. (2022).
@@ 2 Related work
Biological neural systems encoding. Brain encoding is concerned with mapping the input stimuli to the neural activations in the brain. Learning this mapping has been heavily investigated in the literature Zhou et al. (2022); Conwell et al. (2021b); Lahner (2022), where most of the approaches conduct a form of deep regression. The study of brain encoding has been advanced by the release of naturalistic neuroscience datasets and benchmarks, with text, audio, image or video stimuli. One of the well established benchmarks that studied how deep networks compare to biological neural systems is the Brain-Score benchmark and framework Schrimpf et al. (2018) which relied on grayscale image stimuli. Another recent well established dataset and benchmark, The Algonauts project Cichy et al. (2019, 2021); Gifford et al. (2023), released datasets and challenges that focused on stimuli as natural objects images Cichy et al. (2019), action videos Cichy et al. (2021); Lahner et al. (2023) and natural scenes Gifford et al. (2023). In these benchmarks, fMRI responses were recorded from different subjects and used to study how the human brain encodes these different kinds of stimuli.
Recent works investigated the ability of deep networks to regress on the brain responses for different stimuli Conwell et al. (2021b); Zhou et al. (2022), where one approach Zhou et al. (2022) focused on action videos from The Algonauts Project 2021. However, they only worked with single-image deep neural network architectures. The regression on the brain responses can be conducted with simple ridge regression to more sophisticated methods that do not assume one-to-one mapping and leave the regression to learn re-weighing of different layers Zhou et al. (2022). In this work, we follow this approach closely, but we focus on studying video understanding models to draw insights on how the brain understands actions and models dynamics. While some works in neuroscience studied the time aspect  Zhuang et al. (2021); Nishimoto et al. (2011); Khosla et al. (2021); Nishimoto (2021); Lahner et al. (2023); Güçlü & Van Gerven (2017); Shi et al. (2018); Sinz et al. (2018); Huang et al. (2023), they did not focus on large-scale comparison to deep video understanding. Our work focuses on studying the state-of-the-art deep video understanding models from a neuroscience perspective, in the first large-scale study of its kind. Moreover, we tackle the important question of whether we can differentiate single image vs. video modelling techniques. A recent work Han et al. (2023) questioned whether the comparison between deep networks and biological neural systems can help us identify the inner workings of the brain with the simple question “if we got it right, would we know”. They designed a real setup where the target model is the brain and demonstrated their results using the Brain-Score framework with recorded responses from primates Majaj et al. (2015); Freeman et al. (2013). More importantly, they proposed a simulated setup where the target model is another deep network representation from the different layers. Nonetheless, their real and simulated experiments focused on single-image architectures and stimuli. We aim to bridge the gap and study whether system identification can be attained when working with video understanding models in both simulated and real environments. Moreover, we provide insights on how deep video understanding models compare to the human visual cortex. Additionally, we study self-supervised video representation learning methods that are not explored in previous works.
Video understanding and dynamics modelling. Video understanding is the task of identifying the objects, actions and events within a video and can include object detection/segmentation, action localization and object/actor localization. Vast literature has focused on the sub-task of action recognition especially in deep learning Kong & Fu (2022). The goal of action recognition is to identify the correct action using an input clip sampled from a video. Approaches in action recognition differ in their spatio-temporal and dynamics modelling, methods that relied on 3D convolution Carreira & Zisserman (2017) or some form of it Tran et al. (2018); Li et al. (2019) were the early directions towards that. There was also a widespread use of two-stream architectures inspired by biological neural systems, some required explicit optical flow Carreira & Zisserman (2017) and others relied on extracting slow vs. fast features through controlling the sampling rate of the input clips to each stream Feichtenhofer et al. (2019). Recently, with the wide adoption of transformers in the computer vision field, video transformers started to see wide adoption as well. One of the earliest, TimeSformer Bertasius et al. (2021), focused on comparing sparse vs. dense and spatial vs. temporal vs. unified spatiotemporal modelling. Another recent development in the transformer-based approaches is the multiscale vision transformer (MViT) Fan et al. (2021). All of the aforementioned models are trained in a fully supervised fashion. The recent advances in self-supervised learning and its connection to foundation models that are deployed in various applications, demonstrates the motivation behind our choice of investigating them. Some methods relied on masked modelling Feichtenhofer et al. (2022), others extended it to multi-modal masked modelling Girdhar et al. (2023). In this work, we focus on studying the aforementioned models and more beyond that to investigate the connection between deep networks and the human visual cortex.
@@ 3 Method
In this section, we describe our environment design for the neural encoding and the simulated one. Then, we discuss the deep regression model and encoding technique, followed by a discussion of the candidate video understanding models.
@@ 3.1 Environment design
In biological neural systems encoding, we aim to study how different stimuli map to the recorded brain responses. It is usually studied within the framework of aligning and comparing deep network architectures and biological neural systems. In this case, the biological neural system is considered the target model, and the candidate deep network architecture, that extracts features from the stimuli to be mapped to the brain responses through deep regression, is considered the source model. In general, computational models of brain responses can be evaluated in two schemes: either using a regression on the brain responses from the deep network features extracted or using a representational similarity analysis. In our case, we focus only on the regression on the brain responses as it has practical benefits to predict responses with different stimuli.
@@
Our first goal is to answer the question: “Can we perform system identification for the underlying modelling scheme?”. To answer that, inspired by previous work Han et al. (2023), we use the features extracted from known architectures as the target, on which we apply our regression, instead of the brain responses as an upper bound. This allows us to identify the dynamics modelling, which serves as a form of ground truth for comparing different models. However, unlike previous work that focused only on studying the architecture (i.e., convolutional vs. transformer-based), we go beyond that to study the dynamics modelling. We use the dynamics modelling to refer to the model’s ability to learn from dynamic information provided in an input clip and/or static information from a single image. Furthermore, we study the identification across families of models when encoding the brain responses by looking into the statistical significance of the difference between these families. Specifically, families are defined based on: (i) the input, whether models learned from single images or videos encouraging them to learn dynamics and motion, (ii) the supervision, whether they are trained in a fully-supervised framework for a certain downstream task or in a self-supervised manner using unlabeled data, and (iii) the architecture, whether the model architecture relies on local convolutional operations or transformer-based global operations. While previous works focused on the architecture aspect, we argue it is even more important to look into whether the model is learning dynamics (e.g., motion) or simply using static information from a single image. Moreover, it is important to understand the impact of the supervision signal used to train the model. In the following, we describe the design details of both the simulated environment, where the target is yet another deep network representation, and the real environment, where the target is the visual cortex responses.
@@
Simulated environment. We select three target models from our candidate models that are detailed in Table 1. These three target models are from both the image/video understanding models and convolutional-/transformer-based models. Specifically, we use I3D ResNet-50 Carreira & Zisserman (2017), ViT-B Dosovitskiy et al. (2021) and MViT-B Fan et al. (2021). I3D ResNet-50 is a convolutional-based models, while ViT-B and MViT-B are transformer-based models. On the other hand, I3D ResNet-50 and MViT-B are video understanding models, while ViT-B is a single image model. For each target, we use the other candidate models as source. The layers representations are extracted from each target model, with the layers selected as detailed in Table 2. A dimensionality reduction is first conducted on the output representations for computational efficiency reasons.
Real environment. In the real set of experiments, our target is the brain responses. In this case, we use the public fMRI dataset from Mini-Algonauts Cichy et al. (2021). We only use the training set since the test data is not provided. We perform cross-validation over four folds throughout all our experiments. The dataset provides fMRI recordings of ten subjects who watched short video clips of three seconds average duration. Each video and voxel in the brain was represented by a single activation value. We use the brain responses from nine regions of interest of the visual cortex, these are across two levels: (i) early and mid-level visual cortex (V1, V2, V3, and V4), and (ii) high-level visual cortex (EBA, FFA, STS, LOC, and PPA). The early and mid-level visual cortex regions are concerned with lower-level features such as orientations and frequencies, while the high-level ones are concerned with semantics in terms of objects, scenes, bodies, and faces.
3.2 Encoding technique
Inspired by the recent work Zhou et al. (2022), we use a layer-weighted region of interest encoding that takes the hierarchical nature of deep networks into consideration. Initially, we pre-process the input features from the different layers of a candidate model through averaging the features on the temporal dimension. This is followed by performing sparse random projection Li et al. (2006) for dimensionality reduction and computational efficiency reasons. Assume input features for layer, $l$, after dimensionality reduction as, $X_{l}\in\mathbb{R}^{C\times 1}$, with $C$ features. We learn the weights of one fully connected layer to provide the predictions of the voxels of one region of interest in the visual cortex as, $\hat{Y}_{l}=W_{l}X_{l}$. Where $W_{l}\in\mathbb{R}^{N\times C}$, $\hat{Y}\in\mathbb{R}^{N\times 1}$ and $N$ is the number of voxels in the region of interest. Instead of simple linear or ridge regression, we learn a weighted sum of the predictions of all layers and use the following loss to train our regression model,
@@ 3.2 Encoding technique
Table 1: List of the candidate models and their families and configurations that were used during their training. We list the backbone/s, the training datasets, and the configuration as frame length $\times$ sampling rate. For the training datasets we use ImageNet Deng et al. (2009) (IN), Kinetics-400 Kay et al. (2017) (K400), Charades Kay et al. (2017) (Ch) and Something-something v2 Sigurdsson et al. (2016) (SSV2).
@@
Image Fully-supervised Convolutional ResNet (R50,18-IN-$8\times 8$)
@@
Table 2: Detailed description per architecture of the number of layers and which ones.
Architecture # Layers Sampled Layers
CSN, R(2+1)D, 3DResNet 6
5 convolutional blocks and last fully connected layer
C2D, I3D 7
5 convolutional blocks, max pooling and last fully connected layer
SlowFast 12
5 convolutional blocks for each branch (slow & fast) and the last 2 layers combined
X3D 6
5 convolutional blocks and last fully connected layer
ViT, TimeSformer 4
Grouped 3 blocks, each block 4 layers and last fully connected layer
Dino - B 4
3 Grouped blocks (4 layers) and CLS token
ResNet 7
5 convolutional blocks, average pooling and last fully connected layer
MViT 5
Grouped 4 blocks (4 layers) and last fully connected layer
OmniMAE, MAE 3
Grouped 3 blocks (4 layers)
stMAE 6
Grouped 6 blocks (4 layers)

### v2

@@ Abstract
Extensive literature has drawn comparisons between recordings of biological neurons in the brain and deep neural networks. This comparative analysis aims to advance and interpret deep neural networks and enhance our understanding of biological neural systems. However, previous works did not consider the time aspect and how the encoding of video and dynamics in deep networks relate to the biological neural systems within a large-scale comparison. Towards this end, we propose the first large-scale study focused on comparing video understanding models with respect to the visual cortex recordings using video stimuli. The study encompasses more than two million regression fits, examining image vs. video understanding, convolutional vs. transformer-based and fully vs. self-supervised models. Our study resulted in both, insights to help better understand deep video understanding models and a novel neural encoding scheme to better encode biological neural systems. We provide key insights on how video understanding models predict visual cortex responses; showing video understanding better than image understanding models, convolutional models are better in the early-mid visual cortical regions than transformer based ones except for multiscale transformers and that two-stream models are better than single stream. Furthermore, we propose a novel neural encoding scheme that is built on top of the best performing video understanding models, while incorporating inter-intra region connectivity across the visual cortex. Our neural encoding leverages the encoded dynamics from video stimuli, through utilizing two-stream networks and multiscale transformers, while taking connectivity priors into consideration. Our results show that merging both intra and inter-region connectivity priors increases the encoding performance over each one of them standalone or no connectivity priors. It also shows the necessity for encoding dynamics to fully benefit from such connectivity priors.
@@ 1 Introduction
There has been a recent increase in studies that compare how deep neural networks process input stimuli to the processing that occurs in the brain whether in humans Zhou et al. (2022); Conwell et al. (2021b); Schrimpf et al. (2018); Cichy et al. (2019; 2021), non-human primates, or rodents Conwell et al. (2021a); Schrimpf et al. (2018). The benefits of these studies are two-fold. First, such a comparison can be used to interpret and have a better understanding of black-box deep neural networks and even provide inspirations on how to improve them. Second, it can provide a better understanding and encoding of biological neural systems. Towards achieving this, recent benchmarks have been released to improve the capabilities of machine learning models in neural encoding and comparing them to biological neural systems Schrimpf et al. (2018); Cichy et al. (2019; 2021); Gifford et al. (2023). The current established benchmark from The Mini-Algonauts Project 2021 Cichy et al. (2021), has provided neuro-imaging data for the brain responses from participants watching short video clips. A further extension of the aforementioned dataset Lahner et al. (2024) provided exhaustive analysis with additional cortical regions for the ventral and dorsal streams. It also provided empirical evidence that temporal information was captured in fMRI recordings of the visual cortex for participants watching video stimuli, through studying the effect of frame shuffling and analyzing the first and last second of these videos. The former showed that shuffling lead to degraded performance confirming that there are dynamics captured in the recorded fMRI, while the later showed that these fMRI recordings encode temporally distinct early and late video snapshots. These developments and benchmarks can enable neuroscientists to study how the brain understands dynamics (e.g., motion), which is critical to study in neural encoding.
Machine learning models have been investigated in encoding such neuro-imaging data Lahner (2022); Zhou et al. (2022); Lahner et al. (2024). However, these studies either focused on a small-scale comparison of few video understanding models or conducted their study on single image based models, neglecting the dynamics aspect. More importantly, previous studies did not conduct a systematic analysis of different model families. To address this gap, we propose the first large-scale comparative study of video understanding in neural encoding that encompasses more than two million regression fits (this refers to source models, their layers, target models, regions, and voxels count). Our study takes various properties into consideration where we study image vs. video understanding, convolutional vs. transformer based, single-stream vs. two-stream and fully supervised vs. self supervised ones. Furthermore, we not only study these video understanding models to encode cortical regions recordings, we also study them in a simulated setup where the target is another neural network, inspired by the single image understanding study Han et al. (2023). Our results show that video understanding models are better than image understanding ones in modelling the human visual cortex recordings. Specifically, two-stream convolutional models and multiscale transformers were the best ones. Interestingly, we are the first to demonstrate the effect of multiscale processing in transformers that improves its ability to capture low level cues (e.g. oriented gradients) consequently improving its performance in the early regions.
Finally, we devise a novel neural encoding mechanism that takes into account encoded dynamics and connectivity priors. Voxel connectivity started to get explored in neural encoding in few recent works Mell et al. (2021); Xiao et al. (2022). However, previous works used this approach in a limited scope in which they studied the connectivity between pairs of regions without considering the connectivity across all regions at once. They also did not consider the combined effect of the local connectivity within the same region (intra-region connectivity) and the global connectivity across the different regions (inter-region connectivity).
In summary, our contributions are three fold:
  • We showcase the first large-scale study of deep video understanding models on two datasets for the human visual cortex where the models include convolutional vs. transformer-based, single vs. two stream and fully vs. self-supervised. Our study is comprehensive with more than 35 models from various families and more than two million regression fits, unlike the recent work Lahner et al. (2024) that showed only three models with limited types.
@@
  • We establish a simulated environment setup using image and video understanding models as the target and a real environment setup with the target model as the human visual cortex.
  • We propose a novel fully integrated encoding model that takes into account intra and inter-region connectivity priors in the visual cortex with features extracted from pre-trained video understanding models. We also show that encoding dynamics is an important aspect to enable the full utilization of such connectivity priors.
@@ 2 Related work
Biological neural systems encoding. Brain encoding is concerned with mapping the input stimuli to the neural activations in the brain. Learning this mapping has been heavily investigated in the literature Zhou et al. (2022); Conwell et al. (2021b); Lahner (2022), where most of the approaches conduct a form of deep regression. The study of brain encoding has been advanced by the release of naturalistic neuroscience datasets and benchmarks, with text, audio, image or video stimuli. One of the well established benchmarks that studied how deep networks compare to biological neural systems is the Brain-Score benchmark and framework Schrimpf et al. (2018) which relied on grayscale image stimuli. Another recent well established dataset and benchmark, The Algonauts project Cichy et al. (2019; 2021); Gifford et al. (2023), released datasets and challenges that focused on stimuli as natural objects images Cichy et al. (2019), action videos Cichy et al. (2021); Lahner et al. (2024) and natural scenes Gifford et al. (2023). In these benchmarks, fMRI responses were recorded from different subjects and used to study how the human brain encodes these different kinds of stimuli. In this work, we focus specifically on recorded fMRI data for participants watching short video clips from Mini-Algonauts 2021 Cichy et al. (2021). Other benchmarks were released with video stimuli focus Popham et al. (2021); Lahner et al. (2024), including the extension of the aforementioned dataset and the most recent benchmark, BOLD moments Lahner et al. (2024). BOLD moments provided an exhaustive analysis and an improved dataset, as such we also evaluate on it as part of our study. Recent works have also investigated the ability of deep networks to regress on the brain responses for different stimuli Conwell et al. (2021b); Zhou et al. (2022), where one approach Zhou et al. (2022) focused on video stimuli. However, they mainly worked with single-image deep neural network architectures. In this work, we follow this approach closely, but we focus on studying video understanding models to draw insights on how the brain understands actions and models dynamics. While some works in neuroscience studied the time aspect  Zhuang et al. (2021); Nishimoto et al. (2011); Khosla et al. (2021); Nishimoto (2021); Lahner et al. (2024); Güçlü & Van Gerven (2017); Shi et al. (2018); Sinz et al. (2018); Huang et al. (2023), they did not focus on large-scale comparison. Our work focuses on the first study of state-of-the-art deep video understanding models from a neuroscience perspective with more than two million regression fits.
Voxel connectivity encoding models. The brain is an interconnected system with local correlations within one region and global correlations across regions  Genç et al. (2016); Li et al. (2022). Few recent works explored the potential of using cortical connectivity in neural encoding models  Mell et al. (2021); Xiao et al. (2022). In Mell et al. (2021), they proposed a model that used predefined source voxels as input to predict a target voxel and compared it to the vanilla stimulus-to-voxels prediction models as the standard neural encoding scheme. Nonetheless, these voxels-to-voxels models are not designed to take stimulus as input and define source voxels in an ad hoc manner. Our work is inspired by that direction, yet we propose a fully integrated model that learns a two-stage architecture, stimulus-to-voxels and voxels-to-voxels. More importantly, our approach takes into consideration all voxels from all visual cortex regions and learns the weighting mechanism, instead of relying on ad hoc non learnable mechanism to define source voxels. Another work Xiao et al. (2022) proposed an encoding approach to improve the neural predictions of high-level visual areas using the predictions of low-level visual areas. On the other hand, our method enables learning from all voxels in the same region and other regions at once within a learnable scheme to leverage inter and intra-region connectivity priors. Additionally, our scheme allows for learning from multiple regions and from the same region in one shot, while previous works mainly used connectivity to one paired region. Finally, we are the first to demonstrate the impact of encoding dynamics on utilizing these connectivity priors to improve neural encoding performance.
@@ 3 Method
In this section, we describe our environment design for the real and the simulated experiments. Then, we discuss the candidate video understanding models and our proposed neural encoding scheme.
@@ 3.1 Environment design
In biological neural systems encoding, we aim to study how different stimuli map to the recorded brain responses. It is usually studied within the framework of aligning and comparing deep network architectures and biological neural systems. In this case, the biological neural system is considered the target model, and the candidate deep network architecture, that extracts features from the stimuli to be mapped to the brain responses through deep regression, is considered the source model. However, our first goal is to answer the question: “Can we perform system identification for the underlying dynamics encoding scheme?”. We use the dynamics encoding to refer to the model’s ability to learn from dynamic information provided in an input clip and encode it within the learned representations. To answer that, inspired by previous work Han et al. (2023), we use the features extracted from known architectures as the target, on which we apply our regression, instead of the brain responses as an upper bound. This allows us to identify the dynamics encoding scheme, which serves as a form of ground truth for comparing different models. However, unlike previous work that focused only on studying the architecture (i.e., convolutional vs. transformer-based), we go beyond that to study whether models process standalone images or learn dynamics from the input video (i.e. image vs. video understanding models).
The second question we aim to study: “How do deep video understanding models families compare to biological neural systems?”. Towards this, we study the identification across families of models when encoding the brain responses. Specifically, families are defined based on: (i) the input, whether models learned from single images or videos encouraging them to learn dynamics and motion, (ii) the supervision, whether they are trained in a fully-supervised framework for a certain downstream task or in a self-supervised manner using unlabeled data, and (iii) the architecture, whether the model architecture relies on local convolutional operations or transformer-based global operations. While previous works focused on the architecture aspect, we argue it is even more important to look into whether the model is learning dynamics (e.g., motion) or simply using static information from a single image. Moreover, it is important to understand the impact of the supervision signal used to train the model. Throughout all the experiments we utilize Mini-Algonauts 2021 dataset Cichy et al. (2021), in addition to providing additional results on the BOLD moments dataset Lahner et al. (2024). In the following, we describe the design details of both the simulated environment, where the target is yet another deep network representation, and the real environment, where the target is the visual cortex responses.
@@
Simulated environment. We select four target models from both the image/video understanding models and convolutional-/transformer-based models. Specifically, we use ResNet-50 He et al. (2016), I3D ResNet-50 Carreira & Zisserman (2017), ViT-B Dosovitskiy et al. (2021) and MViT-B Fan et al. (2021b). ResNet-50 and I3D ResNet-50 are convolutional-based models, while ViT-B and MViT-B are transformer-based models. On the other hand, I3D ResNet-50 and MViT-B are video understanding models, while ResNet-50 and ViT-B are single image understanding models. For each target, we use the other candidate models as source. The layers representations are extracted from each target model as detailed in Appendix A, with input videos from Mini-Algonauts. A dimensionality reduction is conducted on these representations for computational efficiency reasons.
@@
Table 1: List of the candidate models and their families and configurations that were used during their training. We list the backbone/s, the training datasets, and the configuration as clip length $\times$ sampling rate. For the training datasets we use ImageNet Deng et al. (2009) (IN), Kinetics-400 Kay et al. (2017) (K400), Charades Sigurdsson et al. (2016) (Ch) and Something-something v2 Goyal et al. (2017) (SSV2).
@@
Image Fully-supervised Convolutional ResNet (R152,101,50,34,18-IN-$8\times 8$)
@@
Real environment. In the real set of experiments, our target is the brain responses. In this case, we use the public fMRI datasets from Mini-Algonauts and BOLD moments Cichy et al. (2021); Lahner et al. (2024). We mainly use the training set, and perform cross-validation over four folds throughout all our experiments. The dataset provides fMRI recordings of ten subjects who watched short video clips of three seconds average duration. Each video and voxel in the brain was represented by a single activation value. In Mini-Algonauts 2021 dataset, we use the brain responses from nine regions of interest of the visual cortex, these are across two levels: (i) early and mid-level visual cortex (V1, V2, V3, and V4), and (ii) high-level visual cortex (EBA, FFA, STS, LOC, and PPA). The early and mid-level visual cortex regions are concerned with lower-level features such as orientations and frequencies, while the high-level ones are concerned with semantics in terms of objects, scenes, bodies, and faces. We also conduct the same study on BOLD moments with a comprehensive 46 cortical regions that is described in the appendix B.1. The datasets we use is provided at TR one second and has shown sensitivity to the temporal ordering of information in the video stimuli Lahner et al. (2024). This motivates our choice for the two datasets when studying video understanding models and their relation to the visual cortex when encoding dynamics.